### v2

@@ 1 Introduction
Typical approaches in economic research assume prior knowledge of many independent clusters. There are many settings where this information is not available, and instead, units in the population have different degrees of connections. [^8] For example, when using villages as clusters, individuals may interact in the same and nearby villages. See Egger et al. (2022) for an empirical example. This paper provides an econometric framework to choose when and how to design the clusters in cluster experiments. Different from existing clustering algorithms geared towards community detection, we motivate the choice of clusters based on the task of estimating global treatment effects. The choice of clustering must balance two competing objectives: the larger the clusters (and the smaller the number of clusters), the smaller the bias of the estimated global effect, but the larger its variance. We introduce an algorithmic procedure – entitled Causal Clustering – to choose the clustering that minimizes a weighted combination of the worst-case bias and variance as a function of the network and clusters. The worst-case approach encodes uncertainty over the dependence of individual outcomes on neighbors’ assignments. We study (i) whether to run a cluster-level instead of individual-level randomization; (ii) how to cluster individuals (and how many clusters to use).
@@
We focus on a class of models where spillover effects are small relative to the outcomes’ variance but possibly non-negligible for inference. This is formalized in a novel framework of local asymptotics where individual outcomes depend arbitrarily on neighbors’ treatments, and, as $n$ grows, spillovers from neighbors (and possibly also direct effects) converge to zero, but at an arbitrary slow rate (e.g., slower than $n^{-1/2}$). This framework encodes the researchers’ uncertainty on the presence (and magnitude) of spillover effects by modeling first-order neighbors’ effect local to zero, with the convergence rate capturing the expected magnitude of spillovers. [^9] The assumption of spillovers within first-order neighbors can be relaxed here by assuming that higher-order spillovers are an order of magnitude smaller than first-order spillovers. The local asymptotic framework we study is consistent with settings with small (but non-negligible) treatment and spillover effects, typical, for instance, in online experiments (e.g., Karrer et al., 2021). We characterize optimal clustering as a function of the expected magnitude of the largest spillover effects that the experiment can generate. The largest size of the spillover effects is a key input in our algorithms, whose characterization, in practice, is necessary for the design of the experiment but can be challenging. This parameter can be informed by previous experiments, in the same spirit of minimum detectable effects used in power analysis (e.g. Baird et al., 2018), or using some particular modeling assumptions. We provide guidance to practitioners in Section 7.
@@
Our analysis proceeds as follows. First, we provide a formal characterization of the worst-case bias and variance. We show that the worst-case bias is closely related to a particular notion of between-clusters connectedness, defined as the per-individual average number of friends in other clusters. The worst-case variance can potentially be an arbitrary function of within-clusters covariances and between-clusters covariances: individuals in the same cluster have identical assignments and in different clusters may share common neighbors. We show that the variance only depends on the average squared cluster size, up to an asymptotically negligible error. This result formalizes the intuition that a larger number of clusters, with a small variance in cluster size, decreases the variance of the estimator.
@@
We draw the implications of these results for choosing between a cluster experiment (for a given clustering) or assigning treatments independently between individuals (i.e., Bernoulli design or completely randomized design). Suppose the magnitude of the spillover effects is smaller than the square root of the number of clusters. In that case, the variance component dominates the bias, and a Bernoulli design is preferred (where a Bernoulli design is a special case of a cluster design with clusters containing a single unit). Vice-versa, a cluster design is preferred if the bias dominates the variance. Intuitively, because our objective trades off the bias and variance of the estimator, whenever the number of clusters is small, it is best to run a Bernoulli design for any value of spillover effects local to zero. On the other hand, if the number of clusters is sufficiently large and the cluster design appropriately controls the bias of the estimator, a cluster design is preferred. We provide practitioners with a simple and explicit decision rule between cluster and Bernoulli designs that can be easily implemented in practice.
@@
We then turn to the design of optimal clustering, where the choice is not whether to run a cluster or Bernoulli design but rather which clustering to use. The choice of optimal clustering reduces to a novel penalized minimum cut optimization problem, with a penalty that depends on the cluster size variation. The program admits a convenient formulation as a sequence of trace-optimization problems, each solved via off-the-shelf semidefinite programming algorithms.
@@
We provide an empirical application using unique network data from the universe of Facebook users. We show that our procedure provides an explicit ranking between different clustering algorithms implemented at scale at Facebook. We also illustrate trade-offs in the choice of the clustering algorithm and properties of the graph. A second application using network data from Cai et al. (2015) illustrates the method’s advantages for choosing clusters in field experiments. We show that the choice of the clusters based on village identity is sub-optimal in this application because the number of village clusters is too small relative to the optimal clustering, and provide an explicit choice of the clustering. These results illustrate the applicability of the method in choosing the clustering and number of clusters.
@@
A particular class of clustering algorithms is $\varepsilon$-net clustering algorithms – which consist of assigning sequentially individuals in the same neighborhood to the same cluster (Ugander et al., 2013; Eckles et al., 2017). Variants of these algorithms have been recently studied in Leung (2022) (and more recently in contemporaneous work by Leung, 2023) for spatial spillovers, and Faridani and Niehaus (2022) in general non-Euclidean spaces. These papers provide an optimal rate of the clusters’ size as a function of how interference decays in space for asymptotically unbiased estimators. Here, instead, we provide an explicit trade-off and novel characterization of the bias and variance, leveraging local asymptotics. It allows us to characterize the optimal clustering as the solution of a novel and simple trace-optimization program (different from the algorithms studied in the previous references). The comparison of a cluster and Bernoulli designs through local asymptotics is also a novel contribution.
@@
Additional references of experiments with networks in the absence of cluster experiments are Basse and Airoldi (2018b), and Jagadeesan et al. (2020) for estimating direct instead of global average treatment effects studied here; Kang and Imbens (2016) who study encouragement designs, without focusing on the problem of mean-squared error optimal designs; Viviano (2020a) for experiments on networks using information from a pilot study; Basse and Airoldi (2018a) discuss limitations of design-based causal inference under interference. None of these papers study (optimal) cluster designs.
@@
The literature on treatment effects under network interference includes Aronow and Samii (2017), Hudgens and Halloran (2008), Manski (2013), Leung (2020), Athey et al. (2018), Goldsmith-Pinkham and Imbens (2013), Sävje et al. (2021), Ogburn et al. (2017), Manresa (2013), Li and Wager (2022), often focusing on first-order degree interference. [^10] Our local asymptotic framework extends to settings with higher order dependence, assuming that higher order neighbors generate spillovers of smaller order compared to first order neighbors, and the network is sufficiently sparse. None of the above references study experimental designs. Finally, we relate more broadly to the literature on clustering in economics – see Wooldridge (2003), Abadie et al. (2017) and references therein – with the difference that here we focus on optimal clustering, rather than taking the clusters as given – and the literature on graph clustering, including Von Luxburg (2007), Newman (2013a, b), Lei (2019), Lei et al. (2020), Li et al. (2022), and references therein. Different from this last set of papers on graph clustering, this paper focuses on treatment effect estimation instead of community detection.
@@ 2 Setup
We consider a setting with $i\in\{1,\cdots,n\}$ units. Let $Y_{i}\in\mathbb{R}$ denote the observed outcome of interest for individual $i$, $D_{i}\in\{0,1\}$ denote a binary treatment assignment, and $\mathbf{D}\in\{0,1\}^{n}$ the vector of treatment assignments of each unit. Define $\mathbf{A}$ a symmetric adjacency matrix with $\mathbf{A}_{i,j}\in\{0,1\}$, and $Y_{i}(\mathbf{d}),\mathbf{d}\in\{0,1\}^{n}$ the potential outcome as a function of the entire vector of treatment assignments, with $Y_{i}=Y_{i}(\mathbf{D})$. We implicitly condition on $\mathbf{A}$, observed to the researchers, and (unobserved) potential outcomes $Y(\mathbf{d})$, unless otherwise specified (see Remark 7 for an extension with unobserved $\mathbf{A}$). We define $\mathcal{N}_{i}=\{j:\mathbf{A}_{i,j}=1\}$ the set of individuals connected to $i$. Let $\mathcal{N}_{n,\max}=\max_{i\in\{1,\cdots,n\}}|\mathcal{N}_{i}|$, the maximum degree. When $|\mathcal{N}_{i}|=0$ for all nodes, we define $\mathcal{N}_{n,\max}=1$ for notational convenience.
Our primary focus is on estimating the global average treatment effect,
@@
The overall effect defines the effect if all units had received the treatment compared to none of the individuals receiving the treatment. This is a target estimand in many economic applications when researchers are interested in implementing the policy on all individuals in the population, e.g., Egger et al. (2022), Muralidharan and Niehaus (2017), Karrer et al. (2021) (see Remark 9 for a discussion on other estimands). Table 4 summarizes the notation.
@@ 2.1 Potential outcomes and spillover effects
For $i\in\{1,\cdots,n\}$,
$Y_{i}(\mathbf{d})=\mu_{i}(\mathbf{d}_{i},\mathbf{d}_{\mathcal{N}_{i}}),\quad\forall\mathbf{d}\in\{0,1\}^{n},$
for some functions $\mu_{i}(1,\cdot)\in\mathcal{M}_{1,i},\mu_{i}(0,\cdot)\in\mathcal{M}_{0,i}$ for some set of functions $\mathcal{M}_{1,i},\mathcal{M}_{0,i}$.
@@
The potential outcomes $\mu=(\mu_{1},\cdots,\mu_{n})\in\mathcal{M}=\otimes_{i=1}^{n}\mathcal{M}_{i}$ where $\mathcal{M}_{i}$ is the set such that $\mu_{i}(1,\cdot)\in\mathcal{M}_{1,i},\mu_{i}(0,\cdot)\in\mathcal{M}_{0,i}$. For all $i\in\{1,\cdots,n\}$, $d\in\{0,1\}$, the function classes $\mathcal{M}_{d,i}$ contains all functions satisfying
@@
    $\mu_{i}(d,\cdot)\in[\psi_{L},\psi_{R}],\forall\mu_{i}(d,\cdot)\in\mathcal{M}_{d,i}$ for some constants $-\infty<\psi_{L}<\psi_{R}<\infty$ that do not vary with $n$;
@@
    for all $\mathbf{d}\in\{0,1\}^{|\mathcal{N}_{i}|}$,
    $\sup_{\mu_{i}(0,\cdot)\in\mathcal{M}_{0,i}}\Big{|}\mu_{i}(0,\mathbf{d})-\mu_{i}(0,\mathbf{0})\Big{|}$ $=\frac{\bar{\phi}_{n}}{|\mathcal{N}_{i}|}\sum_{k\in\mathcal{N}_{i}}\mathbf{d}_{k},$ (2)
    $\sup_{\mu_{i}(1,\cdot)\in\mathcal{M}_{1,i}}\Big{|}\mu_{i}(1,\mathbf{d})-\mu_{i}(1,\mathbf{1})\Big{|}$ $=\frac{\bar{\phi}_{n}}{|\mathcal{N}_{i}|}\sum_{k\in\mathcal{N}_{i}}\Big{(}1-\mathbf{d}_{k}\Big{)},$
    for some $\bar{\phi}_{n}\leq\psi_{R}-\psi_{L}$.
@@
Assumption 2 imposes two restrictions. Condition (i) states that potential outcomes are uniformly bounded. [^11] This restriction is common in the literature, e.g., Kitagawa and Wang (2021), and can be relaxed by assuming random sub-gaussian potential outcomes. Condition (ii) is our main restriction on the exposure mapping. Condition (ii) is attained if $\mathcal{M}_{1,i},\mathcal{M}_{0,i}$ are Lipschitz function classes in the share of treated neighbors. Condition (ii) states that potential outcomes vary in the share of neighbors’ treatments by at most $\bar{\phi}_{n}$. Here $\bar{\phi}_{n}$ captures the magnitude of (largest) spillovers. The condition that $\bar{\phi}_{n}\leq\psi_{R}-\psi_{L}$ is a regularity assumption that guarantees that $(ii)$ is consistent with the boundedness restriction in $(i)$. In Section 6.1, we relax Assumption 2, by imposing $(ii)$ only on the expected value of the potential outcomes (instead of the realized potential outcomes). In Section 6.2, we relax Assumption 2 by allowing for heterogeneity in the range of potential outcomes and the spillover effects.
The component $\bar{\phi}_{n}$ depends on $n$ and will play an important role in our asymptotic analysis as $n\rightarrow\infty$. We focus on settings where $\bar{\phi}_{n}$ is small (i.e., $\bar{\phi}_{n}=o(1)$ as $n\rightarrow\infty$), in the spirit of a local asymptotic framework (e.g. Hirano and Porter, 2009), but its convergence rate can be arbitrarily slow. These scenarios formalize the idea that spillover effects (and possibly but not necessarily also overall treatment effects $\tau_{n}$) are local to zero. For example, in an information campaign, we might expect spillover (and direct) effects to be small but non-negligible for inference. Small treatment and spillover effects are common in many experimental setups. In experiments conducted on online platforms, individual (and spillover) effects are often statistically small, whereas the sum of all such effects can be economically significant with a large target population. Our asymptotic framework allows for such scenario, where $\bar{\phi}_{n}$, while local to zero, can be of order slower than $n^{-1/2}$. Empirical relevant examples include Athey et al. (2023), Karrer et al. (2021), among others.
We show how different magnitudes of spillover effects justify different designs.
Assumption 3 (Local and sparse network asymptotics).
@@
We consider asymptotic scenarios with a sequence of $\Big{(}\mathbf{A},(Y_{i}(\cdot))_{i=1}^{n}\Big{)}$, indexed by $n$, where $\bar{\phi}_{n}\min\{n,\mathcal{N}_{n,\max}^{2}\}=o(1)$.
Assumption 3 formalizes the local asymptotic framework considered here: spillovers converge to zero, but the rate can be arbitrarily slow up to a term that depends on the squared maximum degree. Assumption 3 holds for arbitrary rates of convergence of the spillover effects $\bar{\phi}_{n}$ for networks with bounded degree (e.g. De Paula et al., 2018, where $\mathcal{N}_{n,\max}$ is bounded), and requires faster rates of $\bar{\phi}_{n}$ (smaller spillover effects), for dense networks.
@@
Consider a class of functions of the form $\mu_{i}(\mathbf{d})=\mu(T_{i}(\mathbf{d}))+\varepsilon_{i}$, with $T_{i}(\mathbf{d})=\Big{[}\mathbf{d}_{i},(1-\mathbf{d}_{i})\times\frac{\sum_{j\neq i}\mathbf{A}_{i,j}\mathbf{d}_{j}}{\sum_{j\neq i}\mathbf{A}_{i,j}},\mathbf{d}_{i}\times\frac{\sum_{j\neq i}\mathbf{A}_{i,j}\mathbf{d}_{j}}{\sum_{j\neq i}\mathbf{A}_{i,j}}\Big{]}$, and $\mu(t)=t^{\top}\beta$ for $\beta\in[-\bar{\Delta}_{n},\bar{\Delta}_{n}]\times[-\bar{\phi}_{n},\bar{\phi}_{n}]^{2},$ for some arbitrary $\bar{\Delta}_{n},\bar{\phi}_{n}$, and $\varepsilon_{i}$ that is not a function of $\mathbf{d}$. Then Assumption 2 holds. ∎
Remark 1 (Weaker conditions on the degree).
Assumption 3 can be sharpened by imposing restrictions on the average second-order degree instead of maximum degree; see Equation (33) in the proof of Lemma C.2. We state this version for expositional convenience. ∎
@@
In the second scenario, suppose that the assumption of first-order effects approximates higher-order effects up to a term of smaller order than first-order effects. Specifically, suppose that $Y_{i}(\mathbf{d})=\mu_{i}(\mathbf{d}_{i},\mathbf{d}_{\mathcal{N}_{i}})+\mathcal{O}(h_{n}),$ for some $h_{n}\rightarrow 0$. Our results hold if $h_{n}=o(1/n)$, capturing the idea that first-order effects $\bar{\phi}_{n}$ are larger than second-order effects. In Appendix B we provide an example and sufficient conditions for the approximation error due to higher order interference being of order $o(1/n)$ in the presence of endogenous peer effects (Manski, 1993; Bramoullé et al., 2009), where the individual outcome depends on other units’ outcomes.
In practice, higher-order effects can be small, leading to underpowered studies, especially when individuals have many friends and first-order effects capture most of the spillovers. This motivates our focus on first-order effects. ∎
Remark 3 (Local asymptotics and direct effect).
Our local asymptotic framework also allows direct treatment effects (and global effects) to be local to zero. Specifically, it is possible that $\tau_{n}=o(1)$ at arbitrary rate, e.g., at the same rate as $\bar{\phi}_{n}$. Therefore, our local asymptotic assumption does not require that spillover effects are local to the global effect (since $\tau_{n}$ can also converge to zero). Instead, our local asymptotics formalizes settings where the noise-to-signal ratio decreases slower than $n^{-1/2}$ (for example, common in experiments on online platforms, Karrer et al. (2021)). ∎

### v3

@@ 1 Introduction
Typical approaches in economic research assume prior knowledge of many independent clusters. There are many settings where this information is not available, and instead, units in the population have different degrees of connections. [^8] For example, when using villages as clusters, individuals may interact in the same and nearby villages. See Egger et al. (2022) for an empirical example. This paper provides an econometric framework to choose when and how to design the clusters in cluster experiments. Different from existing clustering algorithms geared towards community detection, we motivate the choice of clusters based on the task of estimating global treatment effects. The choice of clustering must balance two competing objectives: the larger the clusters (and the smaller the number of clusters), the smaller the bias of the estimated global effect, but the larger its variance. We introduce an algorithmic procedure – entitled Causal Clustering – to choose the clustering that minimizes a weighted combination of the worst-case bias and variance as a function of the network and clusters. The worst-case approach encodes uncertainty over the dependence of individual outcomes on neighbors’ assignments. We study the optimal clustering (and how many clusters to use) in this scenario.
@@
We focus on a class of models where spillover effects are local (within neighbors) and small relative to the outcomes’ variance but possibly non-negligible for inference. This framework encodes the researchers’ uncertainty on the magnitude of spillover effects by modeling first-order neighbors’ effect (and possibly also direct and overall effects) local to zero possibly at arbitrary rates. [^9] The assumption of spillovers within first-order neighbors can be relaxed here by assuming that higher-order spillovers are an order of magnitude smaller than first-order spillovers. From a practical perspective, we therefore focus on a class of designs which are (approximately) optimal when spillover (and direct) effects are small, common, for instance, in online experiments (e.g., Karrer et al., 2021). We characterize optimal clustering as a function of the expected magnitude of the largest spillover effects that the experiment can generate. The largest size of the spillover effects is a key input in our algorithms, whose characterization, in practice, is necessary for the design of the experiment but can be challenging. This parameter can be informed by previous experiments, in the same spirit of minimum detectable effects used in power analysis (e.g. Baird et al., 2018), or using some particular modeling assumptions. We provide a practical discussion in Section 7.
@@
Our analysis proceeds as follows. First, we provide a formal characterization of the worst-case bias and variance. The worst-case bias is closely related to a particular notion of between-clusters connectedness, defined as the per-individual average number of friends in other clusters, similar for instance to notions of cluster purity (e.g. Brennan et al., 2022; Toulis and Kao, 2013). On the other hand, the worst-case variance can potentially be an arbitrary function of within-clusters covariances and between-clusters covariances: individuals in the same cluster have identical assignments and in different clusters may share common neighbors. We show that the variance only depends on the average squared cluster size, up to a small approximation error. This result formalizes the intuition that a larger number of clusters, with a small variance in cluster size, decreases the variance of the estimator. Taken together, these two results provide, as we show, a simple and intuitive expression of the worst-case mean squared error associated with a given clustering.
@@
We draw the implications of these results for choosing between a cluster experiment (for a given clustering) or assigning treatments independently between individuals (i.e., Bernoulli design or completely randomized design). Suppose the magnitude of the spillover effects is smaller than the square root of the number of clusters. In that case, the variance component dominates the bias, and a Bernoulli design is preferred (where a Bernoulli design is a special case of a cluster design with clusters containing a single unit). Vice-versa, a cluster design is preferred if the bias dominates the variance. We provide practitioners with a simple and explicit decision rule between cluster and Bernoulli designs that can be easily implemented.
@@
We then turn to the design of optimal clustering, i.e., which clustering and how many clusters to use. The choice of optimal clustering reduces to a novel sequence of penalized minimum cut optimization problems, with a penalty that depends both on a measure of cluster purity and (different from standard clustering algorithm) on the variance contribution of each cluster. Each problem is solved via off-the-shelf semidefinite programming algorithms.
@@
We provide an empirical application using unique network data from the universe of Facebook users. We show that our procedure provides an explicit ranking between different clustering algorithms implemented at scale at Facebook. We also illustrate trade-offs in the choice of the clustering algorithm and properties of the graph. A second application using network data from Cai et al. (2015) illustrates the method’s advantages for choosing clusters in field experiments. We show that the choice of the clusters based on village identity is sub-optimal in this application because the number of village clusters is too small relative to the optimal clustering because the number of village clusters is too small relative to the optimal clustering, and provide an explicit choice of the clustering, and provide an explicit choice of the clustering. These results illustrate the applicability of the method in choosing the clustering and number of clusters.
@@
A particular class of clustering algorithms is $\varepsilon$-net clustering algorithms – which consist of assigning sequentially individuals in the same neighborhood to the same cluster (Ugander et al., 2013; Eckles et al., 2017). Variants of these algorithms have been recently studied in Leung (2022b) (and more recently in contemporaneous work by Leung, 2023) for spatial spillovers, and Faridani and Niehaus (2022) in general non-Euclidean spaces. These papers provide an optimal rate of the clusters’ size as a function of how interference decays in space for asymptotically unbiased estimators. Here, instead, we provide an explicit trade-off and novel characterization of the bias and variance, leveraging local asymptotics.
In particular, as Leung (2022b) conjectures, network dependence is difficult to handle for general cluster experiments. A key challenge is that the estimator’s variance depends not only on the network structure but also on the clustering structure: outcomes in two different clusters are dependent if they share common neighbors or if they are in the cluster with some cross-connection between the two clusters. The lack of explicit and simple-to-optimize expressions for the variance have motivated designs that either focus on minimizing the bias only (e.g. Brennan et al., 2022; Taylor and Eckles, 2018), or on rate-optimal designs only (Leung, 2022b; Faridani and Niehaus, 2022). Our local asymptotic framework allows us to explicitly characterize the variance through a simple and intuitive expression. This result allows us to characterize the optimal clustering as the solution of a novel and simple trace-optimization program (different from the algorithms studied in the previous references).
@@
Additional references of experiments with and without networks in the absence of cluster experiments are Basse and Airoldi (2018b), and Jagadeesan et al. (2020) for estimating direct instead of global average treatment effects studied here; Kang and Imbens (2016) who study encouragement designs, without focusing on the problem of mean-squared error optimal designs; Viviano (2020a) for experiments on networks using information from a pilot study; Basse and Airoldi (2018a) discuss limitations of design-based causal inference under interference. Recent econometric literature studying experimental design focusing however on $i.i.d.$ data include Cai and Rafi (2024); Tabord-Meehan (2023); Armstrong (2022) among others. None of these papers study (optimal) cluster designs.
@@
The literature on treatment effects under network interference includes Aronow and Samii (2017), Hudgens and Halloran (2008), Toulis and Kao (2013), Manski (2013), Leung (2020), Athey et al. (2018), Goldsmith-Pinkham and Imbens (2013), Sävje et al. (2021), Ogburn et al. (2017), Manresa (2013), Li and Wager (2022), often focusing on first-order degree interference. [^10] Our local asymptotic framework extends to settings with higher order dependence, assuming that higher order neighbors generate spillovers of smaller order compared to first order neighbors, and the network is sufficiently sparse. None of the above references study experimental designs. Finally, we relate more broadly to the literature on clustering in economics – see Wooldridge (2003), Abadie et al. (2017) and references therein – with the difference that here we focus on optimal clustering, rather than taking the clusters as given – and the literature on graph clustering, including Von Luxburg (2007), Newman (2013a, b), Lei (2019), Lei et al. (2020), Li et al. (2022), and references therein. Different from this last set of papers on graph clustering, this paper focuses on treatment effect estimation instead of community detection.
@@ 2 Setup
We consider a setting with $i\in\{1,\cdots,n\}$ units. Let $Y_{i}\in\mathbb{R}$ denote the observed outcome of interest for individual $i$, $D_{i}\in\{0,1\}$ denote a binary treatment assignment, and $\mathbf{D}\in\{0,1\}^{n}$ the vector of treatment assignments of each unit. Define $\mathbf{A}$ a symmetric adjacency matrix with $\mathbf{A}_{i,j}\in\{0,1\}$, and $Y_{i}(\mathbf{d}),\mathbf{d}\in\{0,1\}^{n}$ the potential outcome as a function of the entire vector of treatment assignments, with $Y_{i}=Y_{i}(\mathbf{D})$. We implicitly condition on $\mathbf{A}$, observed to the researchers, and (unobserved) potential outcomes $Y(\mathbf{d})$, unless otherwise specified (see Remark 7 for an extension with unobserved $\mathbf{A}$). We define $\mathcal{N}_{i}=\{j:\mathbf{A}_{i,j}=1\}$ the set of individuals connected to $i$. To avoid dividing by zero, we define $|\mathcal{N}_{i}|=\max\{1,\sum_{j}\mathbf{A}_{i,j}\}$. Let $\mathcal{N}_{n,\max}=\max_{i\in\{1,\cdots,n\}}|\mathcal{N}_{i}|$, the maximum degree. We focus on estimating the global average effect,
@@
The overall effect defines the effect if all units had received the treatment compared to none of the individuals receiving the treatment. This is a target estimand in many applications when researchers are interested in implementing the policy at scale, e.g., Egger et al. (2022), Muralidharan and Niehaus (2017), Karrer et al. (2021) (see Appendix B.4 for a discussion on other estimands). Table 2 summarizes the notation.
@@ 2.1 Potential outcomes and spillover effects
For $i\in\{1,\cdots,n\}$, $Y_{i}(\mathbf{d})=\mu_{i}(\mathbf{d}_{i},\mathbf{d}_{\mathcal{N}_{i}}),\forall\mathbf{d}\in\{0,1\}^{n},$ for some functions $\mu_{i}(1,\cdot)\in\mathcal{M}_{1,i},\mu_{i}(0,\cdot)\in\mathcal{M}_{0,i}$ for some set of functions $\mathcal{M}_{1,i},\mathcal{M}_{0,i}$.
@@
The potential outcomes $\mu=(\mu_{1},\cdots,\mu_{n})\in\mathcal{M}=\otimes_{i=1}^{n}\mathcal{M}_{i}$ where $\mathcal{M}_{i}$ is the set such that $\mu_{i}(1,\cdot)\in\mathcal{M}_{1,i},\mu_{i}(0,\cdot)\in\mathcal{M}_{0,i}$. For all $i\in\{1,\cdots,n\}$, $d\in\{0,1\}$, the function class $\mathcal{M}_{d,i}$ is the class of functions that include all functions $\mu_{i}(d,\cdot)$ satisfying the following two conditions:
@@
    $\mu_{i}(d,\cdot)\in[\psi_{L},\psi_{R}]$ for some constants $-\infty<\psi_{L}<\psi_{R}<\infty$ that do not vary with $n$;
@@
    for all $\mathbf{d}\in\{0,1\}^{|\mathcal{N}_{i}|}$, for some $\bar{\phi}_{n}\leq\psi_{R}-\psi_{L}$
    $\Big{|}\mu_{i}(0,\mathbf{d})-\mu_{i}(0,\mathbf{0})\Big{|}$ $\leq\frac{\bar{\phi}_{n}}{|\mathcal{N}_{i}|}\sum_{k\in\mathcal{N}_{i}}\mathbf{d}_{k},$ (2)
    $\Big{|}\mu_{i}(1,\mathbf{d})-\mu_{i}(1,\mathbf{1})\Big{|}$ $\leq\frac{\bar{\phi}_{n}}{|\mathcal{N}_{i}|}\sum_{k\in\mathcal{N}_{i}}\Big{(}1-\mathbf{d}_{k}\Big{)}.$
@@
Assumption 2 imposes two restrictions. Condition (i) states that potential outcomes are uniformly bounded. [^11] This restriction is common in the literature, e.g., Kitagawa and Wang (2021), and can be relaxed by assuming random sub-gaussian potential outcomes. Condition (ii) is our main restriction on the exposure mapping. Condition (ii) is attained if $\mathcal{M}_{1,i},\mathcal{M}_{0,i}$ are Lipschitz function classes in the share of treated neighbors. Condition (ii) states that potential outcomes vary in the share of neighbors’ treatments by at most $\bar{\phi}_{n}$. Here $\bar{\phi}_{n}$ captures the magnitude of (largest) spillovers. The condition that $\bar{\phi}_{n}\leq\psi_{R}-\psi_{L}$ is a regularity assumption that guarantees that $(ii)$ is consistent with the boundedness restriction in $(i)$. In Appendix B.1, we relax Assumption 2, by imposing $(ii)$ only on the expected value of the potential outcomes (instead of the realized potential outcomes). In Appendix B.2, we relax Assumption 2 by allowing for heterogeneity in the range of potential outcomes and the spillover effects. [^12] Also, note that because $\mathcal{M}_{d,i}$ contains all functions satisfying $(i)$ and $(ii)$ it also contains constant functions as well as linear functions in the share of treated neighbors with linear coefficient bounded by $\bar{\phi}_{n}$.
The component $\bar{\phi}_{n}$ quantifies the largest magnitude of spillover effects and will play an important role in our asymptotic analysis as $n\rightarrow\infty$. We quantify the magnitude of spillover effects through the following asymptotic framework.
Assumption 3 (Spillover local asymptotics).
@@
We consider a sequence of $\Big{(}\mathbf{A},(Y_{i}(\cdot))_{i=1}^{n}\Big{)}$, indexed by $n$, where $\lim\sup_{n}\bar{\phi}_{n}\min\{n,\mathcal{N}_{n,\max}^{2}\}\leq\varepsilon$, for some finite constant $\varepsilon\geq 0$.
Assumption 3 formalizes the local asymptotic framework considered here, where $\bar{\phi}_{n}\mathcal{N}_{n,\max}^{2}$ denotes the product of the largest spillover effects with the squared maximum degree. [^13] Assumption 3 can be sharpened by imposing restrictions on the average second-order degree instead of maximum degree; see Equation (38) in the proof of Lemma D.2. We assume that such a product is bounded in the limit by $\varepsilon\geq 0$. For $\varepsilon=0$, spillover effects converge to zero asymptotically. This may occur if $\bar{\phi}_{n}\rightarrow 0$ at an arbitrary rate for networks with bounded degree (e.g. De Paula et al., 2018, where $\mathcal{N}_{n,\max}$ is bounded). In this case, the rate of convergence of $\bar{\phi}_{n}$ characterizes the magnitude of the spillover effects. As in a local asymptotic framework (e.g. Hirano and Porter, 2009), rates of converges of order equal or slower than the effective number of observations (allowed here) indicates non-negligible spillover effects even with $\varepsilon=o(1)$. Although we impose no restrictions on $\varepsilon$, we will think of $\varepsilon$ as small. Specifically, we provide a framework that characterizes approximately optimal experimental design, where the quality of the approximation depends on $\varepsilon$. Also, note that overall effects may also exhibit the same (or similar) asymptotic behavior, see Remark 1.
Scenarios where $\bar{\phi}_{n}$ is local to zero formalize the idea that spillover effects (and possibly but not necessarily also overall treatment effects $\tau_{n}$) are small but possibly non-negligible for inference. For example, in contexts with online advertising Lewis and Rao (2015) observe that a treatment effect of 0.005 of a standard deviation of the outcome would be highly profitable for the company given the cost of advertising (see also Athey et al., 2023, for a discussion). Similar examples in digital advertisements with small effects are Athey et al. (2023) with $i.i.d.$ data and Karrer et al. (2021) with network (Facebook) data, where total effect size (sum of direct and spillover effects) are also of the order of 0.005 standard deviations. Motivated by these applications we therefore focus on settings where first-order effects are small but possibly non-negligible. As we show below, this framework allows us to provide explicit characterization of the bias and variance of the clustering, different from models of decaying interference as for example in Leung (2022a), where such characterization is not possible. In other scenarios where either the network is dense, or spillover effects have large magnitudes, the quality of the approximations we provide below will deteriorate.
@@
Consider a class of functions of the form $\mu_{i}(\mathbf{d})=\mu(T_{i}(\mathbf{d}))+\nu_{i}$, with $T_{i}(\mathbf{d})=\Big{[}\mathbf{d}_{i},(1-\mathbf{d}_{i})\times\frac{\sum_{j\neq i}\mathbf{A}_{i,j}\mathbf{d}_{j}}{\sum_{j\neq i}\mathbf{A}_{i,j}},\mathbf{d}_{i}\times\frac{\sum_{j\neq i}\mathbf{A}_{i,j}\mathbf{d}_{j}}{\sum_{j\neq i}\mathbf{A}_{i,j}}\Big{]}$, and $\mu(t)=t^{\top}\beta$ for $\beta\in[-\bar{\Delta}_{n},\bar{\Delta}_{n}]\times[-\bar{\phi}_{n},\bar{\phi}_{n}]^{2},$ for some arbitrary $\bar{\Delta}_{n},\bar{\phi}_{n}$, and $\nu_{i}$ that is not a function of $\mathbf{d}$. Then Assumption 2 holds. ∎
Remark 1 (Direct and overall effect).
Our asymptotic framework also allows direct treatment effects (and global effects) to be of the same magnitude of spillover effects. Specifically, it is possible that $\tau_{n}$ is of the same order of $\bar{\phi}_{n}$. Therefore, our local asymptotic assumption does not require that spillover effects are local to the global effect (since $\tau_{n}$ can also converge to zero). Instead, our local asymptotics formalizes settings where the noise-to-signal ratio can be small (common in experiments on online platforms, Karrer et al. (2021)). ∎
@@
In the second scenario, suppose that the assumption of first-order effects approximates higher-order effects up to a term of smaller order than first-order effects. In this case our results can extend as we discuss in Appendix B.5.
In practice, higher-order effects can be small, leading to underpowered studies, especially when individuals have many friends and first-order effects capture most of the spillovers. ∎